OpenStudy (anonymous):

Simplify 2x + x. I feel like this would be the right answer : 2x²

OpenStudy (anonymous):

that would be right if they had a multiplication however they have addition

OpenStudy (anonymous):

when there's addition combine like terms

OpenStudy (anonymous):

3x ²

OpenStudy (anonymous):

only 3x

OpenStudy (anonymous):

3x² if I am not mistaken

OpenStudy (anonymous):

when combining like terms don't change the power

OpenStudy (anonymous):

explain your answer to me if can.

OpenStudy (debbieg):

2 apples + 1 apple = 3 apples

OpenStudy (anonymous):

thanks so it would be 3x?

OpenStudy (anonymous):

\[x \times x^{3} = x^{4}\] yes

OpenStudy (debbieg):

if x is an apple, then \(x^2\) is an orange. Don't mix apples and oranges! :) Another way to understand this is to use factoring: 2x + x = x(2 + 1) = x(3) = 3x You can't get \(x^2\) unless you have another FACTOR of x. \(x\cdot x = x^2\) but x + x does not.

OpenStudy (anonymous):

when combining like terms just count how often you have some x don't change the power so we have 2x and another x that makes 3x

OpenStudy (debbieg):

In general: \(ax + bx = (a + b)x\) \(ax^2 + bx^2 = (a + b)x^2\) \(ax^3 + bx^3 = (a + b)x^3\) ......etc........... For addition/subtraction, just combine the coefficients of the LIKE TERMS. (apples and apples :)

OpenStudy (anonymous):

thanks a lot! I understand now. Let me write all of this down now.